OpenStudy (anonymous):

A popular phone manufacturer has released a new version of its smart phone. This new version is priced based on the size of its memory card. The data in the table gives a correlation coefficient of 0.96. If the values in the memory column were expressed in MB instead of GB and the values in the price column were expressed in euros instead of dollars, what would the value of the correlation coefficient be? 0.77 0.84 0.96 0.99

OpenStudy (anonymous):

I would think it's just a matter of unit conversion: \[\frac{\text{megabytes}}{\text{dollar}}\times\frac{1\text{ dollar}}{0.74\text{ euro}}\times\frac{1\text{ gigabyte}}{1024\text{ megabytes}}\]

OpenStudy (anonymous):

Slight correction: \[\frac{\text{gigabytes}}{\text{dollar}}\times\frac{1\text{ dollar}}{0.74\text{ euro}}\times\frac{1024\text{ megabytes}}{1\text{ gigabyte}}\]

OpenStudy (anonymous):

so what would be my answer

OpenStudy (aum):

I am not quite sure but I think correlation coefficient is a statistical measure of how strongly two sets of data are dependent on each other and so I am inclined to guess it will be independent of the units used.

OpenStudy (anonymous):

^That might also be the case. I'm not sure if there's a unit associated with the coefficient to begin with.

OpenStudy (anonymous):

soo you think it would be a

OpenStudy (aum):

I am guessing the correlation coefficient will remain the same at 0.96.

OpenStudy (anonymous):

i got it right thanks